what is the equation of the line y=
for m=7/8,(5,-5)

Answer by lwsshak3(11628)

what is the equation of the line y=
for m=7/8,(5,-5)
...
m=7/8,(5,-5)
Standard form for equation of a straight line: y=mx+b, m=slope, b=y-intercept
For given problem:
y=7x/8+b
solving for b using given point a (5,-5)
-5=7*5/8+b
-40=35+8b
8b=-75
b=-75/8
Equation of line:
y=7x/8-75/8